OpenStudy (anonymous):

exponent ^6 radical 50 with another exponent 6 equals?

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\sqrt{x^2}=x\)

OpenStudy (anonymous):

\(\sqrt[6]{x^6}=x\)

OpenStudy (ranga):

\[\Large \sqrt[n]{x^m} = x^{m/n}; \quad \sqrt[6]{50^6} = 50^{6/6} = 50\]